OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 1 = 3, what do you know about the solution(s) to this system? Part 2: Solve the following system and show all of your work. x − 2y = 14 x + 3y = 9

OpenStudy (anonymous):

This is what I put. Is it correct? Part 1: 1=3 is a false statement which indicates no solution. Part 2: x − 2y = 14 x + 3y = 9 x - 2y = 14 + 2y + 2y x = 14 + 2y (14 + 2y) + 3y = 9 14 + 2y + 3y = 9 14 + 5y = 9 -14 -14 5y = -5 y = -5 x + 3(-5)= 9 x + -15 = 9 + 15 + 15 x = 24 Solution: x=24, y=-5

OpenStudy (anonymous):

Did you just answer your own question, or do you still need help?

OpenStudy (anonymous):

I want to know if it's correct.

OpenStudy (anonymous):

part 1 is right u must review the part 2 must be a mistake

OpenStudy (anonymous):

And yes, anything is obviously true is all real numbers, anything not true is no solution. Ok i'll check it.

OpenStudy (anonymous):

see this part 5y = -5 y = -5

OpenStudy (anonymous):

Yeah.

Directrix (directrix):

These summary charts may help you review.